# A Review of Large Language Models and Autonomous Agents in Chemistry

### v1

@@ 3.4 Synthesis Prediction and Encoder-decoder Mol-LLMs
Table 3: Encoder-decoder scientific LLMs. aa:“Model Size” is reported as the number of parameters. $b$: The total number of parameters was not reported.
@@
MOLGEN276 $b$ ZINC15 BART Molecule Generation 2024.03
BioT5+157 252M ZINC20, UniRef50, 33M PubMed articles, 339K mol-text pairs from PubChem, 569K FASTA-text pairs from Swiss-prot T5 Molecule Captioning, Molecule Generation, Property Prediction, 2024-02
BioT5172 252M ZINC20, UniRef50, full-articles from BioRxiv and PubMed, mol-text-IUPAC information from PubChem T5 Molecule Captioning, Property Prediction 2023-12
nach0141 250M MoleculeNet, USPTO, ZINC T5 Property prediction, Molecule generation, Question answering, NER 2023.11
ReactionT5174 220M ZINC and ORD T5 Property prediction and Reaction prediction 2023.11
Text+Chem T5158 60M, 220M 11.5M or 33.5M samples curated from Vaucher et al. 277, Toniato et al. 214, and CheBI-20 T5 Molecule Captioning, Product Prediction, Retrosynthesis, Molecule Generation 2023.01
MolT5159 60M, 770M C4 dataset T5 Molecule Captioning and Molecule Generation 2022.11
T5Chem175 220M USPTO T5 Product Prediction, Retrosynthesis, Property Prediction 2022.03
ChemFormer139 45M, 230M 100M SMILES from ZINC-15 BART Product Prediction, Property Prediction, Molecular Generation 2022.01
SMILES transformer183 $b$ ChEMBL24 Transformer Property prediction 2019.11
Molecular Transformer222 12M USPTO Transformer Product prediction 2019.08
The encoder-decoder architecture is designed for tasks involving the translation of one sequence into another, making it ideal for predicting chemical reaction outcomes or generating synthesis pathways from given reactants. We begin with a background on optimal synthesis prediction and describe how earlier machine learning has approached this challenge. Following that, we explain how LLMs have enhanced chemical synthesis prediction and optimization.
@@ 3.4.1 Synthesis Prediction
Once a molecule has been identified through property-directed inverse design, the next challenge is to predict its optimal synthesis, including yield. Shenvi 278 describe how the demanding and elegant syntheses of natural products has contributed greatly to organic chemistry. However, in the past 20 years, the focus has shifted away from complex natural product synthesis towards developing reactions applicable for a broader range of compounds, especially in reaction catalysis.278 Yet, complex synthesis is becoming relevant again as it can be digitally encoded, mined by LLMs, and applied to new challenges. Unlike property prediction, reaction prediction is challenging due to the involvement of multiple molecules. Modifying one reactant requires adjusting all others, with different synthesis mechanisms or conditions likely involved. Higher-level challenges exist for catalytic reactions and complex natural product synthesis. Synthesis can be approached in two ways. Forward synthesis involves building complex target molecules from simple, readily available substances, planning the steps progressively. Retrosynthesis, introduced by E.J. Corey in 1988,279 is more common. It involves working backward from the target molecule, breaking it into smaller fragments whose re-connection is most effective. Chemists choose small, inexpensive, and readily available starting materials to achieve the greatest yield and cost-effectiveness. For example, the first total synthesis of discodermolide280 involved 36 such steps, a 24-step longest linear sequence, and a 3.2% yield. There are many possible combinations for the total synthesis of a target molecule, and the synthetic chemist must choose the most sensible approach based on their expertise and knowledge. However, this approach to total synthesis takes many years. LLMs can now transform synthesis such that structure-activity relationship predictions can be coupled in lock-step with molecule selection based on easier synthetic routes. This third challenge of predicting the optimal synthesis can also lead to the creation of innovative, non-natural compounds, chosen because of such an easier predicted sytnthesis but for which the properties are still predicted to meet the needs of the application. Thus, these three challenges introduced above are interconnected.
@@ 3.4.2 Encoder-decoder mol-LLMs
Before we focus on transformer use, some description is provided on the evolution from RNN and Gated Recurrent Unit (GRU) approaches in concert with the move from template-based to semi-template-based to template-free models. Nam and Kim 281 pioneered forward synthesis prediction using a GRU-based translation model, while Liu et al. 282 reported retro-synthesis prediction with a Long Short-Term Memory (LSTM) based seq2seq model incorporating an attention mechanism, achieving 37.4% accuracy on the USPTO-50K dataset.283 The reported accuracies of these early models highlighted the challenges of synthesis prediction, particularly retrosynthesis. Schneider et al. 283 further advanced retrosynthesis by assigning reaction roles to reagents and reactants based on the product.
@@
Evolving from RNNs and GRUs, the field progressed with the introduction of template-based models. Segler and Waller 284 identified that computational “rule-based systems” often failed because they ignored the molecular context, resulting in “reactivity conflicts.” They prioritized suitable transformation rules describing how atoms and bonds change during reactions, applied in reverse for retrosynthesis. They trained a model on 3.5 million reactions, achieving 95% top-10 accuracy in retrosynthesis and 97% for reaction prediction on a validation set of nearly 1 million reactions from the Reaxys database (1771-2015). Though not transformer-based, their work paved the way for LLMs in synthesis applications. However, template-based models rely on explicit reaction templates from known reactions, which limits their ability to predict novel reactions and they require manual updates for learning new data.
@@
Semi-template-based models offered a balance between rigid template-based methods and flexible template-free approaches. They used interpolation or extrapolation within template-defined spaces to predict a wider range of reactions and to adjust based on new data. In 2021, Somnath et al. 285 introduced a graph-based approach recognizing that precursor molecule topology is largely unchanged during reactions. Their model broke the product molecule into “synthons” and added relevant leaving groups, making results more interpretable.286 Training on the USPTO-50k dataset, 283 they achieved a top-1 accuracy of 53.7%, outperforming previous methods.
@@
It is, however, the template-free approaches that align well with transformer-based learning approaches because they learn retrosynthetic rules from raw training data. This provides significant flexibility and generalizability across various types of chemistry. Template-free models are not constrained by template libraries and so can uncover novel synthetic routes that are undocumented or not obvious from existing reaction templates. To pave the way for transformer use in synthesis, Cadeddu et al. 287 drew an analogy between fragments in a compound and words in a sentence due to their similar rank distributions. Schwaller et al. 288 further advanced this with an LSTM network augmented by an attention-mechanism-based encoder-decoder architecture, using the USPTO dataset.283 They introduced a commonly used “regular expression” (or “regex”) for tokenizing molecules, framing synthesis (or retrosynthesis) predictions as translation problems with a data-driven, template-free sequence-to-sequence model. They tracked which starting materials were actual reactants, distinguishing them from other reagents like solvents or catalysts, and used the regex to uniquely tokenize recurring reagents, as their atoms were not mapped to products in the core reaction.
@@
In 2019, Schwaller et al. 222 first applied a transformer for synthesis prediction, framing the task as translating reactants and reagents into the final product. Their model inferred correlations between chemical motifs in reactants, reagents, and products in the dataset (USPTO-MIT,289 USPTO-LEF, 290 USPTO-STEREO288). It required no handcrafted rules and accurately predicted subtle chemical transformations, outperforming all prior algorithms on a common benchmark dataset. The model handled inputs without a reactant-reagent split, following their previous work,288 and accounted for stereochemistry, making it valuable for universal application. Then, in 2020, for automated retrosynthesis, Schwaller et al. 291 developed an advanced Molecular Transformer model with a hyper-graph exploration strategy. The model set a standard for predicting reactants and other entities, evaluated with four new metrics: coverage, class diversity, round-trip accuracy, and Jensen–Shannon divergence. Constructed dynamically, the hypergraph allowed for efficient expansion based on Bayesian-like probability scores, showing high performance despite training data limitations. Notably, accuracy was improved after the re-synthesis of the target product from the newly generated molecular precursors was included in the model. This round-trip accuracy concept was also used by Chen and Jung 292 and Westerlund et al. 293. Also in 2020, Zheng et al. 294 developed a “template-free self-corrected retrosynthesis predictor” (SCROP) using transformer networks and a neural network-based syntax corrector, achieving 59.0% accuracy on a benchmark dataset.295, 283 This approach outperformed other deep learning methods by over 2% and template-based methods by over 6%.
@@
We now highlight advancements in synthesis prediction using the BART Encoder-Decoder architecture, including Chemformer by Irwin et al. 139. This paper emphasized the computational expense of training transformers on SMILES and the importance of pretraining for efficiency. It showed that models pretrained on task-specific datasets or using only the encoder stack were limited for sequence-to-sequence tasks. After transfer learning, Chemformer achieved state-of-the-art results in both sequence-to-sequence synthesis tasks and discriminative tasks, such as optimizing molecular structures for specific properties. They studied the effects of small changes on molecular properties using pairs of molecules from the ChEMBL database133 with a single structural modification. Chemformer’s performance was tested on the ESOL, Lipophilicity, and Free Solvation datasets.55 Irwin et al. 139 also described their use of an in-house property prediction model, but when models train on calculated data for ease of access and uniformity, they abstract away from real-world chemical properties. We again emphasize the importance of incorporating experimentally derived data into Chemistry LLM research to create more robust and relevant models. Continuously curating new, relevant datasets that better represent real-world chemical complexities will enhance the applicability and transferability of these models.
@@
In 2023, Toniato et al. 214 also applied LLMs to single-step retrosynthesis as a translation problem, but increased retrosynthesis prediction diversity by adding classification tokens, or “prompt tokens,” to the target molecule’s language representation, guiding the model towards different disconnection strategies. Increased prediction diversity has high value by providing out-of-the-box synthetic strategies to complement the human chemist’s work. To measure retrosynthesis accuracy, Li et al. 296 introduced Retro-BLEU, a metric adapted from the BLEU score used in machine translation.297 Despite progress in computer-assisted synthesis planning (CASP), not all generated routes are chemically feasible due to steps like protection and deprotection needed for product formation. Widely accepted NLP metrics like BLEU297 and ROUGE298 focus on precision and recall by computing n-gram overlaps between generated and reference texts. Similarly, in retrosynthesis, reactant-product pairs can be treated as overlapping bigrams. Retro-BLEU uses a modified BLEU score, emphasizing precision over recall, as there is no absolute best route for retrosynthesis. Although not yet applied to LLM-based predictions, this approach has value by allowing future performance comparison with a single standard.
@@
Finally, by expanding the use of encoder-decoder architectures outside synthesis prediction into molecular generation, Fang et al. 276 introduced MOLGEN, a BART-based pretrained molecular language model, in a 2023 preprint updated in 2024. MOLGEN addressed three key challenges: generating valid SMILES strings, avoiding an observed bias that existed against natural product-like molecules, and preventing hallucinations of molecules that didn’t retain the intended properties. Pretrained on 100 million molecules using SELFIES127 and a masked language model approach, MOLGEN predicts missing tokens to internalize chemical grammar. An additional highlight of this work is how MOLGEN uses “domain-agnostic molecular prefix tuning.” This technique integrates domain knowledge directly into the model’s attention mechanisms by adding molecule-specific prefixes, trained simultaneously with the main model across various molecular domains. The model’s parameters would thus be adjusted to better capture the complexities and diversities of molecular structures, and domain-specific insights would be seamlessly integrated. To prevent molecular hallucinations, MOLGEN employs a chemical feedback mechanism, to autonomously evaluate generated molecules for appropriate properties, to guide learning and optimization. Such feedback foreshadows a core aspect of autonomous agents, which is their capacity for reflection. We will explore this further below.
@@ 3.5 Multi-Modal LLMs
Multitask or multimodal methods applied to chemistry are now described. In 2021, Edwards et al. 299 proposed Text2Mol, which retrieved molecules using natural language descriptions as queries. This task required integrating both molecular and natural language modalities, making it a cross-lingual retrieval problem due to the distinct grammar of molecular representations. The researchers built a paired dataset of molecules with corresponding text descriptions, and developed an aligned semantic embedding space for retrieval. This was enhanced with a cross-modal attention-based model for explainability and reranking. One stated aim was to improve retrieval metrics, which would further advance the ability for machines to learn from chemical literature. In their follow-up work in 2022, Edwards et al. 159 continued using SMILES string representations along with text descriptions. The two key applications were first, to generate captions for molecules from SMILES strings and second, to output molecules based on textual descriptions of desired properties. With this in mind, several important challenges remain. First, molecules can have many possible descriptions based on their diverse properties. For example, aspirin can be described by its pain-relieving effects, its use in preventing heart attacks, its chemical structure (an ester and a carboxylic acid connected to a benzene ring in ortho geometry), or its degradation into salicylic acid and ethanoic acid in moist conditions 300. Describing such a molecule therefore requires expertise in different chemistry domains, depending on its functions and properties. In contrast, image captioning for common objects like cats and dogs needs much less domain expertise. Thus, obtaining a large, paired dataset of chemical representations and textual descriptions is challenging. Another issue is that standard evaluation metrics, like BLEU, are inadequate for molecule-language tasks. To address this, Edwards et al. 159 developed MolT5, which pretrained a model on a large corpus of unlabeled natural language text and molecule strings using a denoising objective. They fine-tuned the model on gold standard annotations and improved the metric based on their prior Text2Mol work 299. MolT5 effectively generated both molecule and caption outputs.
@@
In other work, Seidl et al. 301 developed CLAMP. Trained on large biochemical datasets, CLAMP adapts to new tasks using separate modules for chemical and language inputs, and predicts biochemical activity for drug discovery. Xu et al. 302 presented BioTranslator, a method for translating user-written text descriptions into non-text biological data, facilitating the discovery of novel cell types and applications in protein function prediction and drug target identification. This method enabled data exploration beyond controlled vocabularies using free text. ChatDrug, by Liu et al. 303, integrates multimodal capabilities through a prompt module, a retrieval and domain feedback module, and a conversation module for systematic drug editing. It identifies and manipulates molecular structures for better interpretability in pharmaceutical research. Christofidellis et al. 158 introduced a multi-domain, multi-task language model capable of handling tasks across chemical and natural language domains without specific pretraining for each. Liu et al. 304 developed MolXPT, a unified model that integrates text and molecules for enhanced molecular property prediction and translation, demonstrating robust zero-shot molecular generation capabilities. Zhang et al. 305 introduced PremuNet, a pretrained multi-representation fusion network that enhances molecular property prediction by integrating various molecular data representations. Liu et al. 306 developed GIT-Mol, integrating graph, image, and text data, significantly improving molecular property prediction and molecule generation validity. Gao et al. 307 advanced targeted molecule generation with DockingGA, combining transformer neural networks with genetic algorithms and docking simulations for optimal molecule generation, utilizing Self-referencing Chemical Structure Strings to represent and optimize molecules. Zhou et al. 308 developed TSMMG, a teacher-student LLM for multi-constraint molecular generation, learning from a large set of text-molecule pairs and generating molecules that meet complex property requirements. Zhu et al. 309 proposed 3M-Diffusion, a multi-modal molecular graph generation method that uses natural language descriptions to generate diverse, novel molecular graphs. Gong et al. 310 introduced TGM-DLM, a diffusion model for text-guided molecule generation that overcomes limitations of autoregressive models in generating precise molecules from textual descriptions. Fang et al. 311 developed MolTC, a multimodal framework that effectively integrates graphical information of molecules for improved prediction of molecular interactions. Soares et al. 312 describe MULTIMODAL-MOLFORMER, a multimodal language model that predicts molecular properties by integrating chemical language representations with physicochemical features. It uses a causal multistage feature selection method to identify physicochemical features directly impacting target properties. These features are combined with molecular embeddings from MOLFORMER187, significantly enhancing prediction accuracy for complex tasks like biodegradability and PFAS toxicity.

### v2

@@ 3.4 Synthesis Prediction and Encoder-decoder Mol-LLMs
Table 3:  Encoder-decoder scientific LLMs. $a$:“Model Size” is reported as the number of parameters. $b$: The total number of parameters was not reported.
@@
MOLGEN295 $b$ ZINC15 BART Molecule Generation 2024.03
BioT5+129 252M ZINC20, UniRef50, 33M PubMed articles, 339K mol-text pairs from PubChem, 569K FASTA-text pairs from Swiss-prot T5 Molecule Captioning, Molecule Generation, Property Prediction, 2024-02
BioT5144 252M ZINC20, UniRef50, full-articles from BioRxiv and PubMed, mol-text-IUPAC information from PubChem T5 Molecule Captioning, Property Prediction 2023-12
nach0128 250M MoleculeNet, USPTO, ZINC T5 Property prediction, Molecule generation, Question answering, NER 2023.11
ReactionT5147 220M ZINC and ORD T5 Property prediction and Reaction prediction 2023.11
Text+Chem T5130 60M, 220M 11.5M or 33.5M samples curated from Vaucher et al. 296, Toniato et al. 225, and CheBI-20 T5 Molecule Captioning, Product Prediction, Retrosynthesis, Molecule Generation 2023.01
MolT5131 60M, 770M C4 dataset T5 Molecule Captioning and Molecule Generation 2022.11
T5Chem148 220M USPTO T5 Product Prediction, Retrosynthesis, Property Prediction 2022.03
ChemFormer146 45M, 230M 100M SMILES from ZINC-15 BART Product Prediction, Property Prediction, Molecular Generation 2022.01
Text2Mol297 $b$ CheBI-20 SciBERT w/ decoder Molecule captioning and conditional molecule generation 2021.
SMILES transformer156 $b$ ChEMBL24 Transformer Property prediction 2019.11
Molecular Transformer233 12M USPTO Transformer Product prediction 2019.08
The encoder-decoder architecture is designed for tasks involving the translation of one sequence into another, making it ideal for predicting chemical reaction outcomes or generating synthesis pathways from given reactants. We begin with a background on optimal synthesis prediction and describe how earlier machine learning has approached this challenge. Following that, we explain how LLMs have enhanced chemical synthesis prediction and optimization. Although, our context below is aptly chosen to be synthesis prediction, other applications exist. For example, SMILES Transformer (ST)156 is worth a mention, historically, because it explored the benefits of unsupervised pretraining to produce continuous, data-driven molecular fingerprints from large SMILES-based datasets.
@@ 3.4.1 Synthesis Prediction
Once a molecule has been identified through property-directed inverse design, the next challenge is to predict its optimal synthesis, including yield. Shenvi 298 describe how the demanding and elegant syntheses of natural products has contributed greatly to organic chemistry. However, in the past 20 years, the focus has shifted away from complex natural product synthesis towards developing reactions applicable for a broader range of compounds, especially in reaction catalysis.298 Yet, complex synthesis is becoming relevant again as it can be digitally encoded, mined by LLMs,299 and applied to new challenges. Unlike property prediction, reaction prediction is challenging due to the involvement of multiple molecules. Modifying one reactant requires adjusting all others, with different synthesis mechanisms or conditions likely involved. Higher-level challenges exist for catalytic reactions and complex natural product synthesis. Synthesis can be approached in two ways. Forward synthesis involves building complex target molecules from simple, readily available substances, planning the steps progressively. Retrosynthesis, introduced by E.J. Corey in 1988,300 is more common. It involves working backward from the target molecule, breaking it into smaller fragments whose re-connection is most effective. Chemists choose small, inexpensive, and readily available starting materials to achieve the greatest yield and cost-effectiveness. For example, the first total synthesis of discodermolide301 involved 36 such steps, a 24-step longest linear sequence, and a 3.2% yield. There are many possible combinations for the total synthesis of a target molecule, and the synthetic chemist must choose the most sensible approach based on their expertise and knowledge. However, this approach to total synthesis takes many years. LLMs can now transform synthesis such that structure-activity relationship predictions can be coupled in lock-step with molecule selection based on easier synthetic routes. This third challenge of predicting the optimal synthesis can also lead to the creation of innovative, non-natural compounds, chosen because of such an easier predicted sytnthesis but for which the properties are still predicted to meet the needs of the application. Thus, these three challenges introduced above are interconnected.
@@ 3.4.2 Encoder-decoder mol-LLMs
Before we focus on transformer use, some description is provided on the evolution from RNN and Gated Recurrent Unit (GRU) approaches in concert with the move from template-based to semi-template-based to template-free models. Nam and Kim 302 pioneered forward synthesis prediction using a GRU-based translation model, while Liu et al. 303 reported retro-synthesis prediction with a Long Short-Term Memory (LSTM) based seq2seq model incorporating an attention mechanism, achieving 37.4% accuracy on the USPTO-50K dataset.304 The reported accuracies of these early models highlighted the challenges of synthesis prediction, particularly retrosynthesis. Schneider et al. 304 further advanced retrosynthesis by assigning reaction roles to reagents and reactants based on the product.
@@
Evolving from RNNs and GRUs, the field progressed with the introduction of template-based models. As a parallel to the separate development of the Chematica305, 306 synthesis mapping tool, Segler and Waller 307 identified that computational “rule-based systems” often failed because they ignored the molecular context, resulting in “reactivity conflicts.” They prioritized suitable transformation rules describing how atoms and bonds change during reactions, applied in reverse for retrosynthesis. They trained a model on 3.5 million reactions, achieving 95% top-10 accuracy in retrosynthesis and 97% for reaction prediction on a validation set of nearly 1 million reactions from the Reaxys database (1771-2015). Though not transformer-based, their work paved the way for LLMs in synthesis applications. However, template-based models rely on explicit reaction templates from known reactions, which limits their ability to predict novel reactions and they require manual updates for learning new data.
@@
Semi-template-based models offered a balance between rigid template-based methods and flexible template-free approaches. They used interpolation or extrapolation within template-defined spaces to predict a wider range of reactions and to adjust based on new data. In 2021, Somnath et al. 308 introduced a graph-based approach recognizing that precursor molecule topology is largely unchanged during reactions. Their model broke the product molecule into “synthons” and added relevant leaving groups, making results more interpretable.309 Training on the USPTO-50k dataset, 304 they achieved a top-1 accuracy of 53.7%, outperforming previous methods.
@@
It is, however, the template-free approaches that align well with transformer-based learning approaches because they learn retrosynthetic rules from raw training data. This provides significant flexibility and generalizability across various types of chemistry. Template-free models are not constrained by template libraries and so can uncover novel synthetic routes that are undocumented or not obvious from existing reaction templates. To pave the way for transformer use in synthesis, Cadeddu et al. 310 drew an analogy between fragments in a compound and words in a sentence due to their similar rank distributions. Schwaller et al. 311 further advanced this with an LSTM network augmented by an attention-mechanism-based encoder-decoder architecture, using the USPTO dataset.304 They introduced a commonly used “regular expression” (or “regex”) for tokenizing molecules, framing synthesis (or retrosynthesis) predictions as translation problems with a data-driven, template-free sequence-to-sequence model. They tracked which starting materials were actual reactants, distinguishing them from other reagents like solvents or catalysts, and used the regex to uniquely tokenize recurring reagents, as their atoms were not mapped to products in the core reaction.
@@
In 2019, going beyond the ”neural machine" work of Nam and Kim 302, Schwaller et al. 233 first applied a transformer for synthesis prediction, framing the task as translating reactants and reagents into the final product. Their model inferred correlations between chemical motifs in reactants, reagents, and products in the dataset (USPTO-MIT,312 USPTO-LEF, 313 USPTO-STEREO311). It required no handcrafted rules and accurately predicted subtle chemical transformations, outperforming all prior algorithms on a common benchmark dataset. The model handled inputs without a reactant-reagent split, following their previous work,311 and accounted for stereochemistry, making it valuable for universal application. Then, in 2020, for automated retrosynthesis, Schwaller et al. 314 developed an advanced Molecular Transformer model with a hyper-graph exploration strategy. The model set a standard for predicting reactants and other entities, evaluated with four new metrics: coverage, class diversity, round-trip accuracy, and Jensen–Shannon divergence. Constructed dynamically, the hypergraph allowed for efficient expansion based on Bayesian-like probability scores, showing high performance despite training data limitations. Notably, accuracy was improved after the re-synthesis of the target product from the newly generated molecular precursors was included in the model. This round-trip accuracy concept was also used by Chen and Jung 315 and Westerlund et al. 316. Also in 2020, Zheng et al. 317 developed a “template-free self-corrected retrosynthesis predictor” (SCROP) using transformer networks and a neural network-based syntax corrector, achieving 59.0% accuracy on a benchmark dataset.318, 304 This approach outperformed other deep learning methods by over 2% and template-based methods by over 6%.
@@
We now highlight advancements in synthesis prediction using the BART Encoder-Decoder architecture, starting with Chemformer by Irwin et al. 146. This paper emphasized the computational expense of training transformers on SMILES and the importance of pretraining for efficiency. It showed that models pretrained on task-specific datasets or using only the encoder stack were limited for sequence-to-sequence tasks. After transfer learning, Chemformer achieved state-of-the-art results in both sequence-to-sequence synthesis tasks and discriminative tasks, such as optimizing molecular structures for specific properties. They studied the effects of small changes on molecular properties using pairs of molecules from the ChEMBL database179 with a single structural modification. Chemformer’s performance was tested on the ESOL, Lipophilicity, and Free Solvation datasets.57 Irwin et al. 146 also described their use of an in-house property prediction model, but when models train on calculated data for ease of access and uniformity, they abstract away from real-world chemical properties. We again emphasize the importance of incorporating experimentally derived data into Chemistry LLM research to create more robust and relevant models. Continuously curating new, relevant datasets that better represent real-world chemical complexities will enhance the applicability and transferability of these models.
@@
In 2023, Toniato et al. 225 also applied LLMs to single-step retrosynthesis as a translation problem, but increased retrosynthesis prediction diversity by adding classification tokens, or “prompt tokens,” to the target molecule’s language representation, guiding the model towards different disconnection strategies. Increased prediction diversity has high value by providing out-of-the-box synthetic strategies to complement the human chemist’s work. To measure retrosynthesis accuracy, Li et al. 319 introduced Retro-BLEU, a metric adapted from the BLEU score used in machine translation.320 Despite progress in computer-assisted synthesis planning (CASP), not all generated routes are chemically feasible due to steps like protection and deprotection needed for product formation. Widely accepted NLP metrics like BLEU320 and ROUGE321 focus on precision and recall by computing n-gram overlaps between generated and reference texts. Similarly, in retrosynthesis, reactant-product pairs can be treated as overlapping bigrams. Retro-BLEU uses a modified BLEU score, emphasizing precision over recall, as there is no absolute best route for retrosynthesis. Although not yet applied to LLM-based predictions, this approach has value by allowing future performance comparison with a single standard.
@@
Finally, by expanding the use of encoder-decoder architectures outside synthesis prediction into molecular generation, Fang et al. 295 introduced MOLGEN, a BART-based pretrained molecular language model, in a 2023 preprint updated in 2024. MOLGEN addressed three key challenges: generating valid SMILES strings, avoiding an observed bias that existed against natural product-like molecules, and preventing hallucinations of molecules that didn’t retain the intended properties. Pretrained on 100 million molecules using SELFIES173 and a masked language model approach, MOLGEN predicts missing tokens to internalize chemical grammar. An additional highlight of this work is how MOLGEN uses “domain-agnostic molecular prefix tuning.” This technique integrates domain knowledge directly into the model’s attention mechanisms by adding molecule-specific prefixes, trained simultaneously with the main model across various molecular domains. The model’s parameters would thus be adjusted to better capture the complexities and diversities of molecular structures, and domain-specific insights would be seamlessly integrated. To prevent molecular hallucinations, MOLGEN employs a chemical feedback mechanism, to autonomously evaluate generated molecules for appropriate properties, to guide learning and optimization. Such feedback foreshadows a core aspect of autonomous agents, which is their capacity for reflection. We will explore this further below.
@@ 3.5 Multi-Modal LLMs
Multitask or multimodal methods applied to chemistry are now described. In 2021, Edwards et al. 297 proposed Text2Mol, which retrieved molecules using natural language descriptions as queries. This task required integrating both molecular and natural language modalities, making it a cross-lingual retrieval problem due to the distinct grammar of molecular representations. The researchers built a paired dataset of molecules with corresponding text descriptions, and developed an aligned semantic embedding space for retrieval. This was enhanced with a cross-modal attention-based model for explainability and reranking. One stated aim was to improve retrieval metrics, which would further advance the ability for machines to learn from chemical literature. In their follow-up work in 2022, Edwards et al. 131 continued using SMILES string representations along with text descriptions. The two key applications were first, to generate captions for molecules from SMILES strings and second, to output molecules based on textual descriptions of desired properties. With this in mind, several important challenges remain. First, molecules can have many possible descriptions based on their diverse properties. For example, aspirin can be described by its pain-relieving effects, its use in preventing heart attacks, its chemical structure (an ester and a carboxylic acid connected to a benzene ring in ortho geometry), or its degradation into salicylic acid and ethanoic acid in moist conditions 322. Describing such a molecule therefore requires expertise in different chemistry domains, depending on its functions and properties. In contrast, image captioning for common objects like cats and dogs needs much less domain expertise. Thus, obtaining a large, paired dataset of chemical representations and textual descriptions is challenging. Another issue is that standard evaluation metrics, like BLEU, are inadequate for molecule-language tasks. To address this, Edwards et al. 131 developed MolT5, which pretrained a model on a large corpus of unlabeled natural language text and molecule strings using a denoising objective. They fine-tuned the model on gold standard annotations and improved the metric based on their prior Text2Mol work 297. MolT5 effectively generated both molecule and caption outputs.
@@
In other work, Seidl et al. 323 developed CLAMP. Trained on large biochemical datasets, CLAMP adapts to new tasks using separate modules for chemical and language inputs, and predicts biochemical activity for drug discovery. Xu et al. 324 presented BioTranslator, a method for translating user-written text descriptions into non-text biological data, facilitating the discovery of novel cell types and applications in protein function prediction and drug target identification. This method enabled data exploration beyond controlled vocabularies using free text. ChatDrug, by Liu et al. 325, integrates multimodal capabilities through a prompt module, a retrieval and domain feedback module, and a conversation module for systematic drug editing. It identifies and manipulates molecular structures for better interpretability in pharmaceutical research. Christofidellis et al. 130 introduced a multi-domain, multi-task language model capable of handling tasks across chemical and natural language domains without specific pretraining for each. Shoghi et al. 326 describe Joint Multi-domain Pre-training (JMP), based on the hypothesis that pre-training on a diverse set of chemical domains should lead to better generalization towards a foundational model. The pre-training task is framed as a multi-task supervised learning problem, where each label of each pre-training dataset is treated as a separate task. Liu et al. 327 developed MolXPT, a unified model that integrates text and molecules for enhanced molecular property prediction and translation, demonstrating robust zero-shot molecular generation capabilities. Zhang et al. 328 introduced PremuNet, a pretrained multi-representation fusion network that enhances molecular property prediction by integrating various molecular data representations. Liu et al. 329 developed GIT-Mol, integrating graph, image, and text data, significantly improving molecular property prediction and molecule generation validity. Gao et al. 330 advanced targeted molecule generation with DockingGA, combining transformer neural networks with genetic algorithms and docking simulations for optimal molecule generation, utilizing Self-referencing Chemical Structure Strings to represent and optimize molecules. Zhou et al. 331 developed TSMMG, a teacher-student LLM for multi-constraint molecular generation, learning from a large set of text-molecule pairs and generating molecules that meet complex property requirements. Zhu et al. 332 proposed 3M-Diffusion, a multi-modal molecular graph generation method that uses natural language descriptions to generate diverse, novel molecular graphs. Gong et al. 333 introduced TGM-DLM, a diffusion model for text-guided molecule generation that overcomes limitations of autoregressive models in generating precise molecules from textual descriptions. Fang et al. 334 developed MolTC, a multimodal framework that effectively integrates graphical information of molecules for improved prediction of molecular interactions. Soares et al. 335 describe MULTIMODAL-MOLFORMER, a multimodal language model that predicts molecular properties by integrating chemical language representations with physicochemical features. It uses a causal multistage feature selection method to identify physicochemical features directly impacting target properties. These features are combined with molecular embeddings from MOLFORMER196, significantly enhancing prediction accuracy for complex tasks like biodegradability and PFAS toxicity.